# Fuzzy dark matter simulations

### v1

@@ 4.2.4 Cosmological simulations
Validation of numerical convergence with increasing resolution can be challenging, and sometimes prohibitively expensive, for FDM cosmological simulations, especially for wave schemes with higher $m_{22}$ or larger simulation volumes. This difficulty arises primarily from the need to (i) use smaller time steps ($\Delta t\propto 1/m_{22}\sigma_{\rm 1D}^{2}$) and (ii) apply higher spatial resolution ($\Delta h\propto 1/m_{22}\sigma_{\rm 1D}$) across a larger fraction of the simulation domain than in CDM simulations. An added complication is the potential formation of spurious halos (see Section 3.3). A practical and efficient alternative for assessing numerical convergence is to perform collisionless $N$-body simulations using the same initial conditions (Liao et al. 2025). First, the halo masses and the density profiles outside the central solitons in the genuine FDM and $N$-body simulations should agree closely. Note that a good fit to an NFW profile beyond the soliton does not guarantee numerical accuracy, as an incorrect profile may still match an NFW model but with erroneous halo mass and concentration. Second, the halo positions should be consistent across the two types of simulations, since the spatial distribution of spurious halos is primarily driven by numerical noise (see Fig. 8).
@@
Figure 16: Density profiles of the central halo in Fig. 15 at $z=0$ obtained with different evolution schemes from the same initial condition. For the FDM simulation using a hybrid scheme with sufficient resolution (solid line), the central core is well fitted by the soliton solution (thin dotted lines), and the outer profile matches the collisionless $N$-body simulation (thick dotted line). When the same hybrid scheme is applied but with adequate wave-scheme resolution only inside the soliton (dashed line), both the central soliton and the outer profile become overly concentrated, although the halo mass remains accurate, as indicated by the corresponding halo virial radius (vertical lines). In contrast, using the global Fourier method with insufficient resolution to resolve the de Broglie wavelength severely underestimates the halo mass (dash-dotted line)
@@
Fig. 16 compares the density profiles of the central halo in Fig. 15 at $z=0$ obtained with different evolution schemes. This halo has a mass $M_{\rm h}=5.9\times 10^{11}\,h^{-1}\rm{M}_{\odot}$ and a virial radius $r_{\rm h}=166\,h^{-1}\rm kpc$. The central soliton has a mass $M_{\rm s}=2.2\times 10^{9}\,h^{-1}\rm{M}_{\odot}$ and a radius $r_{\rm s}=1.1\,h^{-1}\rm kpc$. The free-fall velocity outside the halo is approximately $110\,\rm km\,\rm s^{-1}$, corresponding to a de Broglie wavelength of $\lambda_{\rm dB,ff}\sim 7.6\,h^{-1}\rm kpc$. For the hybrid scheme with fiducial resolution, the central profile aligns well with the soliton solution (Eq. 23), and the outer profile closely matches that of a collisionless $N$-body simulation using the GADGET-2 code with the same initial condition. In contrast, the global Fourier method, performed with a $1024^{3}$ grid and $3.9\,h^{-1}\rm kpc$ spatial resolution, underestimates the halo mass because the resolution is insufficient to properly resolve $\lambda_{\rm dB,ff}$. When the same hybrid scheme is applied with adequate resolution confined to the soliton but inadequate resolution elsewhere in the halo, quantum pressure and turbulence are underestimated, leading to unphysical halo contraction. The deepened gravitational potential raises the halo temperature and, in turn, increases the soliton energy and mass. Nevertheless, note that (i) the resulting central profile still fits a soliton solution, and (ii) the halo mass remains in good agreement with the fiducial value. The latter is attributed to the fluid scheme employed outside the halo. These observations suggest that agreement in halo mass and a good match to the theoretical soliton profile are necessary but not sufficient to establish numerical accuracy (Liao et al. 2025).
@@
Due to the extremely high computational cost, most FDM cosmological simulations to date are limited to probing relatively small FDM particle masses ($m_{22}\sim 0.1\text{--}2$)—either in small simulation boxes ($L\sim 1\text{--}2\,h^{-1}\rm Mpc$) evolved to $z\sim 0\text{--}1$, or in somewhat larger boxes ($L\sim 5\text{--}20\,h^{-1}\rm Mpc$) but only to higher redshifts ($z\gtrsim 3$). Representative examples include simulations using the global Fourier method, applied to either pure FDM scenarios (Woo and Chiueh 2009; Li et al. 2019; May and Springel 2021, 2022; Chan et al. 2022) or a mixed CDM–FDM model (Laguë et al. 2024), and the wave-based finite-difference scheme with AMR (Schive et al. 2014a, b). Simulations employing hybrid schemes and zoom-in techniques enable resolving more massive halos, simulating larger volumes, or reaching lower redshifts (Veltmaat et al. 2018; Schwabe and Niemeyer 2022; Kunkel et al. 2025; Chan et al. 2025; Chiu et al. 2025). In comparison, $N$-body (Schive et al. 2016; Sarkar et al. 2016; Iršič et al. 2017; Corasaniti et al. 2017; Armengaud et al. 2017; Kobayashi et al. 2017; Ni et al. 2019; Leong et al. 2019; Nadler et al. 2025) and SPH (Zhang et al. 2018a; Nori et al. 2019; Nori and Baldi 2021; Nori et al. 2023) simulations can explore much higher $m_{22}$ and larger volumes at lower redshifts, albeit each with its own limitations. A few pioneering simulations have incorporated baryonic physics (Mocz et al. 2019, 2020; Veltmaat et al. 2020), though they are currently restricted to $m_{22}=2.5$, $L\sim 2\,h^{-1}\rm Mpc$, and $z\sim 4\text{--}5$.
@@ 5 Conclusions
Several alternative approaches are also worth noting. Eigenmode methods offer efficient tools for constructing FDM halos or simulating restricted halo regions, provided that gravitational backreaction from non-dark-matter components can be neglected. Collisionless $N$-body methods, while neglecting quantum pressure, remain valuable for probing large-scale structure and assessing the numerical convergence of genuine FDM cosmological simulations. Finally, machine learning, though not addressed in this review, offers an interesting direction for accelerating or augmenting FDM simulations (e.g., Mishra and Tolley 2025).

### v2

@@ 4.2.4 Cosmological simulations
Validation of numerical convergence with increasing resolution can be challenging, and sometimes prohibitively expensive, for FDM cosmological simulations, especially for wave schemes with higher $m_{22}$ or larger simulation volumes. This difficulty arises primarily from the need to (i) use smaller time steps ($\Delta t\propto 1/m_{22}\sigma_{\rm 1D}^{2}$) and (ii) apply higher spatial resolution ($\Delta h\propto 1/m_{22}\sigma_{\rm 1D}$) across a larger fraction of the simulation domain than in CDM simulations. An added complication is the potential formation of spurious halos (see Section 3.3). A practical and efficient alternative for assessing numerical convergence is to perform collisionless NN-body simulations using the same initial conditions (Liao et al. 2025). First, the halo masses and the density profiles outside the central solitons in the genuine FDM and NN-body simulations should agree closely. Note that a good fit to an NFW profile beyond the soliton does not guarantee numerical accuracy, as an incorrect profile may still match an NFW model but with erroneous halo mass and concentration. Second, the halo positions should be consistent across the two types of simulations, since the spatial distribution of spurious halos is primarily driven by numerical noise (see Fig. 8).
@@
Figure 16: Density profiles of the central halo in Fig. 15 at $z=0$ obtained with different evolution schemes from the same initial condition. For the FDM simulation using a hybrid scheme with sufficient resolution (solid line), the central core is well fitted by the soliton solution (thin dotted lines), and the outer profile matches the collisionless NN-body simulation (thick dotted line). When the same hybrid scheme is applied but with adequate wave-scheme resolution only inside the soliton (dashed line), both the central soliton and the outer profile become overly concentrated, although the halo mass remains accurate, as indicated by the corresponding halo virial radius (vertical lines). In contrast, using the global Fourier method with insufficient resolution to resolve the de Broglie wavelength severely underestimates the halo mass (dash-dotted line)
@@
Fig. 16 compares the density profiles of the central halo in Fig. 15 at $z=0$ obtained with different evolution schemes. This halo has a mass $M_{\rm h}=5.9\times 10^{11}\,h^{-1}\rm{M}_{\odot}$ and a virial radius $r_{\rm h}=166\,h^{-1}\rm kpc$. The central soliton has a mass $M_{\rm s}=2.2\times 10^{9}\,h^{-1}\rm{M}_{\odot}$ and a radius $r_{\rm s}=1.1\,h^{-1}\rm kpc$. The free-fall velocity outside the halo is approximately $110\,\rm km\,\rm s^{-1}$, corresponding to a de Broglie wavelength of $\lambda_{\rm dB,ff}\sim 7.6\,h^{-1}\rm kpc$. For the hybrid scheme with fiducial resolution, the central profile aligns well with the soliton solution (Eq. 23), and the outer profile closely matches that of a collisionless NN-body simulation using the GADGET-2 code with the same initial condition. In contrast, the global Fourier method, performed with a $1024^{3}$ grid and $3.9\,h^{-1}\rm kpc$ spatial resolution, underestimates the halo mass because the resolution is insufficient to properly resolve $\lambda_{\rm dB,ff}$. When the same hybrid scheme is applied with adequate resolution confined to the soliton but inadequate resolution elsewhere in the halo, quantum pressure and turbulence are underestimated, leading to unphysical halo contraction. The deepened gravitational potential raises the halo temperature and, in turn, increases the soliton energy and mass. Nevertheless, note that (i) the resulting central profile still fits a soliton solution, and (ii) the halo mass remains in good agreement with the fiducial value. The latter is attributed to the fluid scheme employed outside the halo. These observations suggest that agreement in halo mass and a good match to the theoretical soliton profile are necessary but not sufficient to establish numerical accuracy (Liao et al. 2025).
@@
Due to the extremely high computational cost, most FDM cosmological simulations to date are limited to probing relatively small FDM particle masses ($m_{22}\sim 0.1\text{--}2$)—either in small simulation boxes ($L\sim 1\text{--}2\,h^{-1}\rm Mpc$) evolved to $z\sim 0\text{--}1$, or in somewhat larger boxes ($L\sim 5\text{--}20\,h^{-1}\rm Mpc$) but only to higher redshifts ($z\gtrsim 3$). Representative examples include simulations using the global Fourier method, applied to either pure FDM scenarios (Woo and Chiueh 2009; Li et al. 2019; May and Springel 2021, 2022; Chan et al. 2022) or a mixed CDM–FDM model (Laguë et al. 2024), and the wave-based finite-difference scheme with AMR (Schive et al. 2014a, b). Simulations employing hybrid schemes and zoom-in techniques enable resolving more massive halos, simulating larger volumes, or reaching lower redshifts (Veltmaat et al. 2018; Schwabe and Niemeyer 2022; Kunkel et al. 2025; Chan et al. 2025; Chiu et al. 2025). In comparison, NN-body (Schive et al. 2016; Sarkar et al. 2016; Iršič et al. 2017; Corasaniti et al. 2017; Armengaud et al. 2017; Kobayashi et al. 2017; Ni et al. 2019; Leong et al. 2019; Nadler et al. 2025) and SPH (Zhang et al. 2018a; Nori et al. 2019; Nori and Baldi 2021; Nori et al. 2023) simulations can explore much higher $m_{22}$ and larger volumes at lower redshifts, albeit each with its own limitations. A few pioneering simulations have incorporated baryonic physics (Mocz et al. 2019, 2020; Veltmaat et al. 2020), though they are currently restricted to $m_{22}=2.5$, $L\sim 2\,h^{-1}\rm Mpc$, and $z\sim 4\text{--}5$.
@@ 5 Conclusions
Several alternative approaches are also worth noting. Eigenmode methods offer efficient tools for constructing FDM halos or simulating restricted halo regions, provided that gravitational backreaction from non-dark-matter components can be neglected. Collisionless NN-body methods, while neglecting quantum pressure, remain valuable for probing large-scale structure and assessing the numerical convergence of genuine FDM cosmological simulations. Finally, machine learning, though not addressed in this review, offers an interesting direction for accelerating or augmenting FDM simulations (e.g., Mishra and Tolley 2025).